A horizontal bar of negligible weight has an 82.6 lb weight hanging from the left end and a 43.0 lb weight hanging from the right end. The bar is seen to balance 56.2 in. from the right end. Find the length of the bar.
Answer by venugopalramana(3286)

A horizontal bar of negligible weight has an 82.6 lb weight hanging from the left end and a 43.0 lb weight hanging from the right end. The bar is seen to balance 56.2 in. from the right end. Find the length of the bar.
LET AB BE THE BAR AND ITS LENGTH BE L.WE HAVE 82.6 POUNDS AT A AND 43 POUNDS AT B.IT IS BALANCING AT C SAY AND WE ARE GIVEN THAT CB=56.2..HENCE TAKING MOMENTS ABOUT C WE GET
82.6*AC=43*CB
82.6(AB-CB)=43*CB
82.6(L-56.2)=43*56.2
82.6L-82.6*56.2=43*56.2
82.6L=43*56.2+82.6*56.2=56.2(43+82.6)=56.2*125.6
L=56.2*125.6/82.6=85.4566 INCHES